Suppose Melody owns a business giving piano lessons. Assume that the market for piano lessons is perfectly competitive and that the equilibrium price of a piano lesson is $20. Melody’s total costs vary depending on the number of piano lessons she offers each day, as shown in the table below.

Number of lessons per day Total cost per day ($)
0 30
1 50
2 68
3 78
4 96
5 115
6 138
7 168
8 208

a. When Melody gives 3 lessons per day, what is her average variable cost?

b. What is the profit-maximizing number of lessons for Melody to give each day?

c. What is Melody's daily economic profit at her profit-maximizing number of lessons?
Answer #1
Number of lessons per day Total cost per day ($) TR FC VC AVC MC PROFIT
0 30 0 30 0 -30
1 50 20 30 20 20 20 -30
2 68 40 30 38 19 18 -28
3 78 60 30 48 16 10 -18
4 96 80 30 66 16.5 18 -16
5 115 100 30 85 17 19 -15
6 138 120 30 108 18 23 -18
7 168 140 30 138 19.71429 30 -28
8 208 160 30 178 22.25 40 -48

TR=P*Q, VC=TC-FC, MC=change in TC/change in Q, Profit = TR-TC, AVC=VC/Q

a) When Melody gives 3 lessons per day, her average variable cost is 16

b) Profit maximizing number of lessons for Melody is where P=MC = 5

c) Daily economic profit = TR-TC = 100-115 = -15